OpenStudy (jasbe):

Find the 2nd term of an arithmetic sequence if t1=0 and t5=16 ? 8 6 4 2

OpenStudy (anonymous):

5-1=4 16/4=4 t1 =0 t2 =0+4 =4 t3 = 4+4 =8 t4 =8+4=12 t5 = 12+4=16

OpenStudy (jasbe):

So @jhonyy9 so the answer is t4 ?

OpenStudy (anonymous):

ask there the second term - yes ? so how many you see that is the second term there ?

OpenStudy (anonymous):

t2 =?

OpenStudy (jasbe):

@jhonyy9 4 ?

OpenStudy (anonymous):

yes with 4 you are right sure

OpenStudy (jasbe):

Ok thank you @jhonyy9

OpenStudy (anonymous):

np was my pleasure good luck bye bye